# Study on Light Propagation through Space-Time Random Media via Stochastic Partial Differential Equations

### v1

@@ Abstract
where $C_{1}(\alpha,H)$ is a specific parameter connected with $H,\alpha$ [61]. Formula Eq.(12) is the asymptotic behavior of the solution when the temporal variable $t$ in the hyperbolic Anderson model tends to infinity. As the space power law index $\alpha$ appears explicitly in the expression Eq.(13), so the spatial properties of the media can be represented in the probability density distribution of light field. Moreover, Eq.(13) gives a theoretical basis for the experimental conclusion in [51, 41, 61].
@@
Eq.(14) and Eq.(15) demonstrate that when the Hurst exponent satisfies $H>1/2$, the random field $W(t,\boldsymbol{x})$ exhibits long-range positive temporal dependence. Under this condition, the light field $u_{w}(t,\boldsymbol{x})$ at any fixed spatial location $\boldsymbol{x}_{0}$ likewise displays persistent temporal correlation. Conversely, for $H<1/2$, a negatively correlated random field induces analogous anti-persistent behavior in the resulting light field. This dual dependence is also corroborated by Eq.(12) and Eq.(13), wherein the identity $(4H-\alpha)/(2-\alpha)\equiv 1$ holds exactly at $H=1/2$, marking the transition between correlation regimes.
@@
Experiment.—To validate the theoretical framework and results proposed in this study, we designed an outdoor experimental setup comprising two functionally integrated systems: a series of impedance-tunable temperature probes to measure the fluctuations of the refractive index square $\mu(t,\boldsymbol{x})$ at two distinct spatial points, and a large-aperture coherent detection system to characterize the amplitude and phase of the propagated light field within the outdoor environment as characterized through measurements of the outdoor random media.
To meet the requirements for experimental validation, key parameters in this design were chosen accordingly. In particular, a laser linewidth of 0.85 kHz was selected to enable precise amplitude and phase measurements over extended distances under realistic conditions, since the phase-stability time of the laser exceeds the fluctuation time of the turbulence-induced random media. An intermediate frequency (IF) of 80 MHz was adopted to ensure sufficient phase resolution within the short-time Fourier transform (STFT) frequency bins. The STFT bin width was set to 10 kHz, which exceeds the Nyquist sampling rate required to capture the characteristic fluctuation scale of the media (approximately 1 kHz) [75], thereby guaranteeing faithful signal reconstruction.
@@
Upon completion of the experimental procedures, we first determined $\mu(t_{0},\boldsymbol{x})$ at a fixed time $t_{0}$ across multiple propagation distances $r$. The power-law scaling behavior was identified through a two-segment piecewise fit in double logarithmic coordinates, minimizing the overall residuals. With the receiver aperture set to 2 mm and the IF signal sampled at 250 MHz, Fig.2 (b1)-(b3) present the relationship between $\dfrac{\langle I^{q}\rangle}{\langle I\rangle^{q}}\bigg|_{\boldsymbol{x}_{0}}$ and $\dfrac{\langle I^{2}\rangle}{\langle I\rangle^{2}}\bigg|_{\boldsymbol{x}_{0}}$ at $q=4,5,6$, confirming the validity of Eq.(13) as proposed by the model. Comparisons with previous experiments [51] and simulations can also be found in the supplementary materials [61].
Subsequently, the receiver aperture was enlarged from 2 mm to 300 mm in order to examine the validity of Eq.(16), with the data acquisition card operating at its maximum sampling rate of 1 GHz. To quantitatively characterize the scaling behavior, the rescaled range analysis (R/S) method [61] was utilized to estimate the Hurst index associated with fluctuations in both the random media and the optical phase [58, 61], as illustrated in Fig.3 (a1)–(a2):
@@
Conclusion.—In this letter, we derive a stochastic partial differential equation that models light propagation in space-time random media. The model incorporates temporal fractional Brownian motion and a random field with spatial power-law correlations to capture perturbation processes in non-Markovian transparent random media. Several theoretical properties of the model are analyzed. Corresponding outdoor experiments were designed and carried out to validate the proposed approach.
The theoretical framework shows direct applicability to critical photonic technologies such as free-space optical communication, coherent imaging through scattering layers, and turbulence-resilient imaging systems. It also holds potential for diagnosing various turbulence regimes via spatial spectral analysis [44], as well as for characterizing types and parameters of random media across practical applications.
Acknowledgments.—This study was carried out under the auspices of.

### v2

@@ Abstract
where $C_{1}(\alpha,H)$ is a specific parameter connected with $H,\alpha$ [62]. Formula Eq.(12) is the asymptotic behavior of the solution when the temporal variable $t$ in the hyperbolic Anderson model tends to infinity. As the space power law index $\alpha$ appears explicitly in the expression Eq.(13), the space-time characteristics of the media are reflected in the distribution of the propagated light field.
@@
Eq.(14) and Eq.(15) demonstrate that when the Hurst exponent satisfies $H>1/2$, the random field $W(t,\boldsymbol{x})$ exhibits long-range positive temporal dependence. Under this condition, the light field $u_{w}(t,\boldsymbol{x})$ at any fixed spatial location $\boldsymbol{x}_{0}$ likewise displays persistent temporal correlation. Conversely, for $H<1/2$, a negatively correlated random field induces analogous anti-persistent behavior in the resulting light field [58, 12]. This dual dependence is also corroborated by Eq.(12) and Eq.(13), wherein the identity $(4H-\alpha)/(2-\alpha)\equiv 1$ holds exactly at $H=1/2$, marking the transition between correlation regimes.
@@
Experiment.—To validate the theoretical framework and results proposed in this study, we designed an outdoor experimental setup comprising two functionally integrated systems: a series of impedance-tunable temperature probes to measure the fluctuations of the refractive index square $\mu(t,\boldsymbol{x})$ at two distinct spatial points, and a large-aperture coherent detection system to characterize the amplitude and phase of the propagated light field within the outdoor environment as characterized by measurements of the outdoor random media.
To meet the requirements for experimental validation, key parameters in this design were chosen accordingly. In particular, a laser linewidth of 0.85 kHz was selected to enable precise amplitude and phase measurements over extended distances under realistic conditions, since the phase-stability time of the laser exceeds the fluctuation time of the turbulence-induced random media. An intermediate frequency (IF) of 80 MHz was adopted to ensure sufficient phase resolution within the short-time Fourier transform (STFT) frequency bins. The STFT bin width was set to 10 kHz, which exceeds the Nyquist sampling rate required to capture the characteristic fluctuation scale of the media (approximately 1 kHz [76]), thereby guaranteeing faithful signal reconstruction.
@@
Following the experiments, we first determined $\mu(t_{0},\boldsymbol{x})$ at a fixed time $t_{0}$ across various propagation distances $r$. The power-law scaling behavior was characterized using a two-segment piecewise fit in double-logarithmic coordinates, minimizing the global residuals. For these measurements, the receiver aperture was set to 2 mm, and the signal after balance detection was sampled at 250 MHz, Fig.2 (b1)-(b3) present the relationship between $\dfrac{\langle I^{q}\rangle}{\langle I\rangle^{q}}\bigg|_{\boldsymbol{x}_{0}}$ and $\dfrac{\langle I^{2}\rangle}{\langle I\rangle^{2}}\bigg|_{\boldsymbol{x}_{0}}$ at $q=4,5,6$, confirming the validity of Eq.(13) as proposed by the model. Comparisons with previous experiments [42, 53] and simulations can also be found in the supplementary materials [62].
Subsequently, the receiver aperture was enlarged from 2 mm to 300 mm in order to examine the validity of Eq.(16), with the data acquisition card operating at its maximum sampling rate of 1 GHz. To quantitatively characterize the scaling behavior, the rescaled range analysis (R/S) method [49, 28, 62] was utilized to estimate the Hurst index associated with fluctuations in both the random media and the phase of light field, as illustrated in Fig.3 (a1)–(a2):
@@
Conclusion.—In this letter, we derive a stochastic partial differential equation that models light propagation in space-time random media. The model incorporates temporal fractional Brownian motion and a random field with spatial power-law correlations to capture perturbation processes in non-Markovian transparent random media. Several theoretical properties of the model are analyzed, and outdoor experiments were conducted to validate the proposed approach. Moreover, this work establishes a theoretical foundation for the experimental findings in [53, 42, 62].
The theoretical framework shows direct applicability to critical photonic technologies such as free-space optical communication, coherent imaging through random transparent media, and turbulence-resilient imaging systems. It also holds potential for diagnosing various turbulence regimes via spatial spectral analysis [45], as well as for characterizing types and parameters of random media across practical applications.
Acknowledgments.—This study was carried out under the auspices of National Key Research and Development Program of China (2024YFF0505602).